OpenStudy (anonymous):

find absolute min and max value of the function g(x)=x^(2/3) find absolute min and max value of the function g(x)=x^(2/3) @Mathematics

OpenStudy (mathmathmath):

find the derivative of the function and set it equal to zero to get your critical points

OpenStudy (anonymous):

thanks, so g'(x)=(2/3)x^(-1/3) right? then (2/3)x^(-1/3)=0, if I multiply each side by 3, do I get 2x(-1/3)=0?, divide by 2, do I get x(-1/3)=0? Then what should I do? Please

OpenStudy (mathmathmath):

x^(2/3) = \[\sqrt[3]{X^2}\]

OpenStudy (anonymous):

okay, how do I find x, if x^(-1/3)=0

OpenStudy (anonymous):

does x=0?

OpenStudy (mathmathmath):

g'(x)=(2/3)x^(-1/3) is correct